Adam is 6 years older but twice Joe's age. Their total age is 27.
What are Adam and Joe's ages?

Answer by richwmiller(17219)

a=6+2j
a+j=27
a=27-j
6+2j=27-j
3j=21
j=7
a=20
check
6+2*7=20
ok